OpenStudy (anonymous):

In a certain country, the official unit of current is the copper leaf. One silver leaf is worth 24 copper leaves, and one gold leaf is worth 62 copper leaves. Lily returns from her vacation to this country and finds that she has 15 more copper leaves than silver leaves, and a total of 31 leaves. If her money is worth 413 copper leaves, how many leaves does she have? a) 19 copper leaves 6 silver leaves 6 gold leaves b) 21 copper leaves 6 silver leaves 4 gold leaves c) 21 copper leaves 4 silver leaves 6 gold leaves d) 19 copper leaves 4 silver leaves 8 gold leaves

OpenStudy (deoxna):

It is a system of equations: \[c = 15 + s\] \[s+c+g=31\] \[62g+24s+c=413\] It is fairly straight forward to solve it from there.. :)